OpenStudy (anonymous):

please help!! square root of 9x minus 8 equals square root of 10

jimthompson5910 (jim_thompson5910):

The problem is \[\large \sqrt{9x-8} = \sqrt{10}\] right?

OpenStudy (anonymous):

yes. Can you help me?

jimthompson5910 (jim_thompson5910):

Square both sides to get \[\large \sqrt{9x-8} = \sqrt{10}\] \[\large (\sqrt{9x-8})^2 = (\sqrt{10})^2\] \[\large 9x-8 = 10\]

jimthompson5910 (jim_thompson5910):

I'm sure you see what to do from here

jimthompson5910 (jim_thompson5910):

Don't forget to check any answers you get

OpenStudy (anonymous):

Yes thank you! So just to be sure all I am basically doing is getting rid or the radical sign than solving for x? The so simple wow.

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

Can you help me with another?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

It looks like this

OpenStudy (anonymous):